### v5

@@ 1 Introduction
The Vehicle Routing Problem (VRP) is a well-known and complex combinatorial problem that has received considerable attention since the fifties. Thus, Dantzig and Ramser [1] formally introduced the Vehicle Routing Problem as a generalization of the Traveling Salesman Problem, where a simple matching-based heuristic for its solution was proposed. A few years later, Clarke and Wright [2] developed one of the most used approaches for this problem with an iterative procedure that enables the rapid selection of an optimal or near-optimal route. Iterative search methods [3, 4, 5] can be adapted easily to problems dealing with a large number of constraints, although they require considerable computing times and several parameter settings. Seminal concepts for exact methods appeared successfully in the eighties when Christofides et. al. [6] proposed tree search algorithms incorporating lower bounds computed from shortest spanning $k$-degree center tree and $q$-paths. Also, Laporte et. al. [7] proposed the Gomory cutting planes and branch and bound approaches based on an integer formulation. Since then, different applications of the VRP have been reported in [8, 9, 10], as well as a variety of exact algorithms [11, 12], heuristics, and metaheuristics [13, 14, 15, 16].
The routing problem treated in this work arises from two well-established variants of the VRP: the Pickup & Delivery Problem (PDP) and the Dial-a-Ride Problem (DARP). These problems have been studied using many approaches, including exact and heuristic methods.
On the one hand, PDP constitutes a relevant subset of routing problems in which commodities or passengers have to be transported from different origins to many other destinations. Savelsbergh and Sol [17] discussed several characteristics distinguishing the PDP from standard VRPs and surveyed these problems and solution methods. Similarly, Parragh et. al. [18] focused on routing problems involving pickups and deliveries related to the transportation of goods from the depot to line-haul customers and from back-haul customers to the depot. Ferrucci [19] considered that clients can be seen as intermediaries who can also return goods. Sartori and Buriol [20] solved the PDP using a matheuristic based on the Iterated Local Search method with an embedded Set Partitioning Problem that is iteratively solved to recombine routes of previously found solutions. Furtado et. al. [21] proposed a modeling strategy that allows the assignment of vehicles to routes explicitly in two-index flow formulations. Extensive surveys on this problem concerning exact and heuristic approaches are available at [22, 23, 24, 25].
On the other hand, the DARP can be considered a generalization of the PDP and the VRP with Time Windows since it consists of designing vehicle routes and schedules for individual persons who specify pickup and delivery requests. The most common application can be found in door-to-door transportation services [26, 27]. Furthermore, Madsen et. al. [28] presented a problem characterized by multiple capacities and multiple objectives describing a system for solving a static Dial-a-Ride routing and scheduling problem with time windows. Braekers et. al. [29] adapted to solve small problem instances of the Multi-Depot Heterogeneous DARP exactly. In [30], a dynamic DARP with time window constraints together with a two-phase insertion algorithm based on route perturbations is considered. If transportation requests of users between a set of pickup points and a set of delivery points are considered, Masson et. al. [31] provided a solution method based on an Adaptive Large Neighborhood Search metaheuristic and explained how to check the feasibility of a request insertion. Complete surveys on DARP can be found in [32, 33, 34].
Joint vehicle and crew routing and scheduling problems are also reported in the literature. Thus, Lam et. al. [35] introduced a problem in which crews can interchange vehicles, resulting in space and time interdependencies between vehicle routes and crew routes. It proposes a constraint programming model that overlays crew routing constraints over a standard VRP. In [36], a project motivated by the Austrian Red Cross is proposed. A solution procedure is provided for the daily planning of home healthcare providers that operate multiple vehicles to deliver nurses to clients and to pick them up after service. Moreover, the authors introduced a matheuristic consisting of two stages, identifying potential walking routes and optimizing the transport system. Molenbruch et. al. [37] introduced a routing algorithm and integrated scheduling procedure to enforce the synchronization of flexible vehicle routes to timetables of the remaining public transport services, enabling the design and operational implementation of an integrated mobility system. In the same context, Drexl et. al. [38] studied a simultaneous vehicle and crew routing and scheduling problem arising in long-distance road transport in Europe; a solution heuristic based on a two-stage decomposition of the problem is developed, considering the European Union social legislation for drivers.
@@
This paper focuses on a version of the joint vehicle and crew routing and scheduling problem. The problem arises from data collection procedures designed at the National Statistics Bureau of Ecuador (INEC) to compute the national Consumer Price Index (CPI). A group of pollsters and a fleet of vehicles traverse shared pedestrian and vehicular routes to visit stores scattered throughout the city of Guayaquil and its outskirts for several days. A pollster visits each store to collect the prices of specific goods needed to calculate the CPI. Furthermore, pollsters can walk or be transported by fixed-capacity vehicles through a workday with morning and afternoon shifts, separated by a mandatory break. Every vehicular route must begin at INEC headquarters, dropping and picking up pollsters and allowing vehicles to return or wait at each store. At the end of each workday, vehicles and pollsters must return to the headquarters. This problem aims to reduce the sum of fixed and variable daily costs. The former is associated with the depot’s maintenance, and the latter is related to hiring pollsters and vehicles.
@@
This paper makes several contributions to the vast literature on Transportation Problems. First, a problem, which is a variation of the General Pickup and Delivery Problem (GPDP) and the Dial-a-Ride Problem (DARP), is proposed. Unlike the GPDP, where transhipments at intermediate locations and autonomous load movements are not allowed, it considers the possibility that a pollster, left at a specific node to perform a task, might be picked up by another vehicle; a pollster can also move through the network from one node to another using a pedestrian route. Thus, transhipments and autonomous transportation of the freight are included in this model. This fact reveals the need for syncing vehicular and pedestrian routes, respecting a maximum arrival time corresponding to the length of a working day. Second, this research studies the problem’s complexity, proving it is $\mathcal{NP}$–hard. The latter result leads to modeling, algorithmic, and computational challenges, which have been adequately overcome using Integer Programming tools and a three-phase heuristic. Finally, the results presented in this study help better understand the pollstering logistics for a public institution and a considerable reduction of operational costs of up to 50%.
@@
The structure of this paper is given as follows: in Section 2, an IP-based formulation for the Integrated Vehicle and Pollster Routing Problem is introduced, along with the proof of its computational complexity. Due to the hardness of the problem, Section 3 presents a three-phase heuristic. Afterwards, a real-world instance and computational results are described in Section 4. Finally, in Section 5, concluding remarks are reported.
@@ 2.1 Notation
Let $n\in\mathbb{N}$ be the number of stores and let $S$ be the set of days of a planning horizon. In this problem, only one depot is considered due to technical requirements. It is assumed that pollsters must depart and arrive in a vehicle on a daily basis. Define $G:=(V,A)$ a multi-graph with node set given by $V:=\{0,\ldots,2n+1\}$ and the arc set $A\subseteq V\times V$. Each store and the depot are associated with two nodes in $V$, see Figure 1. Here, the depot node is 00, and its duplicate is node $12n+$ 1. Similarly, the subsets $C_{-}:=\{1,\dots,n\}$ and $C_{+}:=\{n+1,\dots,2n\}$ correspond to the stores and their duplicates, respectively. Node duplication allows the modeling of vehicle arrival, service, waiting, and departure times separately. Furthermore, define $C:=C_{-}\cup C_{+}$ which will be collectively addressed as the set of stores.
@@
Figure 1: Node duplication: The depot, labelled 00, and two stores, labelled $i$ and $j$ are associated with their duplicates which are labelled $12n+$ 1, $i+n$, and $j+n$, respectively.
@@
The set $A:=A_{S}\cup A_{W}\cup A_{V}$ contains three different kinds of arcs. The first group $A_{S}:=\big{\{}(i,i+n):i\in C_{-}\big{\}}$ is the set of arcs which connect every store with its duplicate; they are called service arcs. The second group of arcs $A_{W}:=\big{\{}(i,j):i\in C_{+},j\in C_{-},i\neq j+n\big{\}}$, called walking arcs, links every pair of stores. Finally, the vehicle transportation arc set, $A_{V}:=\big{\{}(i,j):i,j\in C,i\neq j,i\neq j+n\big{\}}\cup\{(0,i):i\in C\big{\}}\cup\big{\{}(i,2n+1):i\in C\big{\}}$ represents potential connections between nodes or depots using vehicles. Moreover, for every arc $a\in A_{S}$, the parameter $t_{a}>0$ is the time spent by a pollster collecting data at a store; for each arc $a\in A_{W}$, $t_{a}>0$ is the pedestrian travel time between stores, and, for $a\in A_{V}$, $\tau_{a}\geq 0$ is the vehicular travel time between two nodes ($\tau_{a}=0$ only if $a\in A_{S}$). A representation of the resulting multigraph is displayed in Figure 2.
@@
Visits are carried out by a set $E$ of pollsters transported by a homogeneous fleet of vehicles $K$ with positive integer capacity $Q$. Each operational day has an associated fixed cost of $\kappa_{0}\geq 0$. Similarly, vehicles and pollsters have a fixed hiring cost of $\kappa_{1}\geq 0$ and $\kappa_{2}\geq 0$, respectively, and they work within a daily time frame $[0,B_{\max}]$. Due to labor regulations, pollsters must take a break of $P\geq 0$ minutes starting in the time interval $[T_{0},T_{1}]\subset[0,B_{\max}-P]$.
@@
A vehicular route is a simple path from node 00 to $12n+$ 1 using exclusively arcs in $A_{V}$. A pedestrian route, also called a walking path, is a simple path from $i\in C_{-}$ to $j\in C_{+}$ consisting of an alternating sequence of arcs in sets $A_{S}$ and $A_{W}$. A feasible service route for a pollster is defined as a simple path from 00 to $12n+$ 1 linking pedestrian routes with segments of vehicular routes. Thus, at every node $i$, a pollster stays $t_{i,i+n}$ units of time to accomplish a task respecting the maximum daily duty length. At last, a feasible daily plan consists of different feasible service routes synced according to the duration of each pedestrian and vehicular route, and any sequence of feasible daily plans forms an aggregated daily plan.
@@
The Integrated Vehicle and Pollster Routing Problem (IVPRP) consists of finding an aggregated daily plan to visit every node $i\in C$ exactly once at minimum cost. This plan must hire at most $|K|$ vehicles of capacity $Q$ and $|E|$ pollsters in no more than $|S|$ days. The IVPRP is $\mathcal{NP}$–hard, as proved in the following theorem.
@@
Proof: The $\mathcal{NP}$–hardness is proved by using a polynomial transformation from the Traveling Salesman Problem (TSP). Let $K_{n+1}$ ($n\geq 3$) a complete graph with set of nodes $\{0,1,\dots,n\}$ with weights $c:E(K_{n+1})\to\mathbb{R}_{+}$. An instance of the IVPRP is constructed as follows: an arbitrary node in $V(K_{n+1})$ is chosen as the depot. Without loss of generality let 00 be such a node. The remaining nodes in $V(K_{n+1})$ correspond to every store of our problem, i.e., $C:=C_{-}\cup C_{+}$ with $C_{-}=V(K_{n+1})\setminus\{0\}$ and the multigraph $D$ is constructed as above. The cost function $c$ in the TSP corresponds to the vehicular travel times $\tau$ for arcs $(i,j)$, $i,j\in C$, $i\neq j$ in $A_{V}$ and 00 otherwise. Pedestrian travel time $t_{a}\gg\sum_{a\in A_{V}}\tau_{a}$ and $t_{a}=0$ for $a\in A_{S}$. Moreover, fix $|E|=|K|=|S|=1$, $Q=1$, $P=0$, $B_{\max}=\infty$, and the break time window is set to $[0,B_{\max})$.
If the Hamiltonian cycle $\{0,i,j,k,\dots,0\}$ is a feasible solution of the TSP with cost $H$, then IVPRP has a feasible solution with the same cost given by the vehicular route $\{0,i,i+n,j,j+n,k,k+n,\dots,2n+1\}$, the pedestrian route $\bigcup_{i\in C_{-}}(i,i+n)$, and both forming a service route which is also a feasible daily plan. Conversely, suppose that the IVPRP has a feasible solution with an aggregated daily plan of length less than $\sum_{a}\tau_{a}$. Since the pedestrian travel times are larger than $\sum_{a}\tau_{a}$, then the pedestrian arcs in $A_{W}$ are never used in that solution. Thus, the feasible service route reveals a Hamiltonian cycle for the TSP with the same cost. ∎
Figure 2: Multigraph construction for the IVPRP. Focusing only on the depot and on two nodes $i,j\in C_{-}$: There are only outgoing arcs from 00 and only incoming arcs to $12n+$ 1. One service arc is defined for $i$ and $j$ separately, two walking arcs connect $\{i,j,i+n,j+n\}$, and 18 vehicle arcs connect $\{0,i,j,i+n,j+n,2n+1\}$.

### v6

@@ 1 Introduction
The design of routes for one or several vehicles on a network, satisfying constraints related to the transportation of goods, people, or information, is known as the Vehicle Routing Problem (VRP). On the other hand, crew routing is the process of designing routes for crews to perform specific activities. Real-world applications usually involve vehicle and crew routing tasks that need to be solved in an integrated manner. Considering that crews may be able to interchange vehicles and manage synchronized times, this makes the problem combinatorially complex and more challenging.
@@
This paper focuses on a version of the joint vehicle and crew routing and scheduling problem. The problem arises from data collection procedures designed by the National Statistics Bureau of Ecuador (INEC) to compute the national Consumer Price Index (CPI). A group of pollsters and a fleet of vehicles traverse shared pedestrian and vehicular routes to visit stores scattered throughout the city of Guayaquil and its outskirts over several days. A pollster visits each store to collect the prices of specific goods needed to calculate the CPI. Furthermore, pollsters can walk or be transported by fixed-capacity vehicles during a workday with morning and afternoon shifts, separated by a mandatory break. Every vehicular route must begin at INEC headquarters, dropping off and picking up pollsters and allowing vehicles to return or wait at each store. At the end of each workday, vehicles and pollsters must return to the headquarters. This problem aims to reduce the sum of fixed and variable daily costs. The former is associated with the maintenance of the depot, while the latter is related to hiring pollsters and vehicles.
1.1 Literature review
The Vehicle Routing Problem (VRP) is a well-known and complex combinatorial problem that has received considerable attention since the 1950s. Thus, Dantzig and Ramser [1] formally introduced the VRP as a generalization of the Traveling Salesman Problem, where a simple matching-based heuristic for its solution was proposed. A few years later, Clarke and Wright [2] developed one of the most widely used approaches for this problem with an iterative procedure that enables the rapid selection of an optimal or near-optimal route. Iterative search methods [3, 4, 5] can be easily adapted to problems with a large number of constraints, although they require considerable computing times and several parameter settings. Seminal concepts for exact methods appeared successfully in the 1980s when Christofides et. al. [6] proposed tree search algorithms incorporating lower bounds computed from shortest spanning $k$-degree center trees and $q$-paths. Additionally, Laporte et. al. [7] proposed Gomory cutting planes and branch-and-bound (B&B) approaches based on an integer formulation. Since then, various applications of the VRP have been reported in [8, 9, 10], as well as a variety of exact algorithms [11, 12], heuristics, and metaheuristics [13, 14, 15, 16].
The routing problem addressed in this work arises from two well-established variants of the VRP: the Pickup & Delivery Problem (PDP) and the Dial-a-Ride Problem (DARP). These problems have been studied using a variety of approaches, including exact and heuristic methods.
On the one hand, the PDP constitutes a relevant subset of routing problems in which commodities or passengers have to be transported from different origins to multiple destinations. Savelsbergh and Sol [17] discussed several characteristics distinguishing the PDP from standard VRPs and surveyed these problems and solution methods. Similarly, Parragh et. al. [18] focused on routing problems involving pickups and deliveries related to the transportation of goods from the depot to line-haul customers and from back-haul customers to the depot. Ferrucci [19] considered scenarios where clients can be seen as intermediaries who can also return goods. Sartori and Buriol [20] solved the PDP using a matheuristic based on the Iterated Local Search method with an embedded Set Partitioning Problem that is iteratively solved to recombine routes of previously found solutions. Furtado et. al. [21] proposed a modeling strategy that explicitly assigns vehicles to routes in two-index flow formulations. Wang et. al. [22] developed a variant of the pickup and delivery problem, called the two-echelon pickup and delivery problem using public transport; they formulated a new mathematical model based on a space-time network and adopted an adaptive large neighborhood search (ALNS) algorithm for the problem. Extensive surveys on this problem concerning exact and heuristic approaches are available in [23, 24, 25, 26].
On the other hand, the DARP can be considered a generalization of the PDP and the VRP with Time Windows, as it consists of designing vehicle routes and schedules for individual persons who specify pickup and delivery requests. The most common application can be found in door-to-door transportation services [27, 28]. Furthermore, Madsen et. al. [29] presented a problem characterized by multiple capacities and multiple objectives, describing a system for solving a static Dial-a-Ride routing and scheduling problem with time windows. Braekers et. al. [30] proposed an exact solution approach for small problem instances of the Multi-Depot Heterogeneous DARP. In [31], a dynamic DARP with time window constraints is addressed, along with a two-phase insertion algorithm based on route perturbations. When transportation requests of users between a set of pickup points and a set of delivery points are considered, Masson et. al. [32] provided a solution method based on an ALNS metaheuristic and explained how to check the feasibility of a request insertion. Cortenbach et. al. [33] introduced a mixed-integer linear programming formulation for the Dial-a-Ride Problem with Meeting Points; two preprocessing steps and three valid inequalities are introduced, along with two versions of the Tabu Search metaheuristic, to approximate the optimal solution in large-scale networks. Comprehensive surveys on the DARP can be found in [34, 35, 36].
Joint vehicle and crew routing and scheduling problems are also reported in the literature. For example, Lam et. al. [37] introduced a problem in which crews can interchange vehicles, resulting in space and time interdependencies between vehicle routes and crew routes. They proposed a constraint programming model that overlays crew routing constraints on a standard VRP. In [38], a project motivated by the Austrian Red Cross is discussed. A solution procedure is provided for the daily planning of home healthcare providers operating multiple vehicles to deliver nurses to clients and pick them up after service. Moreover, the authors introduced a matheuristic consisting of two stages: identifying potential walking routes and optimizing the transport system. Molenbruch et. al. [39] introduced a routing algorithm and integrated scheduling procedure to enforce the synchronization of flexible vehicle routes with timetables of public transport services, enabling the design and operational implementation of an integrated mobility system. In the same context, Drexl et. al. [40] studied a simultaneous vehicle and crew routing and scheduling problem arising in long-distance road transport in Europe. A solution heuristic based on a two-stage decomposition of the problem was developed, considering the European Union social legislation for drivers. Lucci et. al. [41] presented a simultaneous vehicle routing and crew scheduling problem (SVRCSP) in which long-distance pickup-and-delivery requests must be fulfilled over a multiday planning horizon, subject to several constraints such as multiple time windows and hour-of-service regulations, among others. The authors developed an algorithm based on the GRASP × ILS metaheuristic. Furthermore, Shen and Li [42] introduced a new model for an integrated electric Vehicle Scheduling Problem (VSP) and Crew Scheduling Problem (CSP). This model employs minimum cost flow formulations for the electric VSP, set partitioning for the CSP, and linking constraints. Due to the nonlinear integer nature of the problem, they proposed a method that combines a matching-based heuristic and an integer linear programming solver.
1.2 Contribution and structure
@@
This paper makes several contributions to the vast literature on Transportation Problems, which is reviewed in the context of this work in the next section. First, a problem, which is a variation of the General Pickup and Delivery Problem (GPDP) and the Dial-a-Ride Problem (DARP), is proposed. Unlike the GPDP, where transshipments at intermediate locations and autonomous load movements are not allowed, this problem considers the possibility that a pollster, left at a specific node to perform a task, might be picked up by another vehicle. Additionally, a pollster can move through the network from one node to another using a pedestrian route. Thus, transshipments and autonomous transportation of the freight are included in this model. This fact highlights the need for syncing vehicular and pedestrian routes while respecting a maximum arrival time corresponding to the length of a working day. Second, this research studies the complexity of the problem, proving it is $\mathcal{NP}$–hard. This result leads to modeling, algorithmic, and computational challenges, which have been addressed using Integer Programming tools and a three-phase heuristic. Finally, the results presented in this study provide a better understanding of pollstering logistics for a public institution and demonstrate a considerable reduction of operational costs of up to 50%.
@@
The structure of this paper is as follows: in Section 2, an IP-based formulation for the Integrated Vehicle and Pollster Routing Problem is introduced, along with a proof of its computational complexity. Due to the hardness of the problem, Section 3 presents a three-phase heuristic. Subsequently, a real-world instance and the computational results are described in Section 4. Finally, concluding remarks are provided in Section 5.
@@ 2.1 Notation
Let $n\in\mathbb{N}$ be the number of stores, and let $S$ be the set of days in the planning horizon. In this problem, only one depot is considered due to technical requirements. It is assumed that pollsters must depart from and return to the depot in a vehicle on a daily basis. Define $G\coloneqq(V,A)$ as a multi-graph with a node set given by $V\coloneqq\{0,\ldots,2n+1\}$ and an arc set $A\subseteq V\times V$. Each store and the depot are associated with two nodes in $V$, as illustrated in Figure 1. Specifically, the depot is represented by node 00, and its duplicate is node $12n+$ 1. Similarly, the subsets $C_{-}\coloneqq\{1,\dots,n\}$ and $C_{+}\coloneqq\{n+1,\dots,2n\}$ correspond to the stores and their duplicates, respectively. Node duplication enables the modeling of vehicle arrival, service, waiting, and departure times separately. Furthermore, define $C\coloneqq C_{-}\cup C_{+}$, which will collectively be referred to as the set of stores.
@@
Figure 1: Node duplication: The depot, labeled 00, and two stores, labeled $i$ and $j$, are associated with their duplicates, labeled $12n+$ 1, $i+n$, and $j+n$, respectively.
@@
The set $A\coloneqq A_{S}\cup A_{W}\cup A_{V}$ contains three different kinds of arcs. The first group, $A_{S}\coloneqq\big{\{}(i,i+n):i\in C_{-}\big{\}}$, is the set of arcs that connect each store with its duplicate; these are called service arcs. The second group of arcs, $A_{W}\coloneqq\big{\{}(i,j):i\in C_{+},j\in C_{-},i\neq j+n\big{\}}$, called walking arcs, links each pair of stores. Finally, the vehicle transportation arc set, $A_{V}\coloneqq\big{\{}(i,j):i,j\in C,i\neq j,i\neq j+n\big{\}}\cup\big{\{}(0,i):i\in C\big{\}}\cup\big{\{}(i,2n+1):i\in C\big{\}}$, represents potential connections between nodes or depots using vehicles. Moreover, for each arc $a\in A_{S}$, the parameter $t_{a}>0$ represents the time spent by a pollster collecting data at a store; for each arc $a\in A_{W}$, $t_{a}>0$ is the pedestrian travel time between stores; and for each arc $a\in A_{V}$, $\tau_{a}\geq 0$ is the vehicular travel time between two nodes ($\tau_{a}=0$ only if $a\in A_{S}$). A representation of the resulting multigraph is displayed in Figure 2.
Figure 2: Multigraph construction for the IVPRP. Focusing only on the depot and two nodes $i,j\in C_{-}$: there are only outgoing arcs from 00 and only incoming arcs to $12n+$ 1. One service arc is defined for $i$ and $j$ separately, two walking arcs connect $\{i,j,i+n,j+n\}$, and 18 vehicle arcs connect $\{0,i,j,i+n,j+n,2n+1\}$.
@@
Visits are carried out by a set $E$ of pollsters transported by a homogeneous fleet of vehicles $K$ with a positive integer capacity $Q$. Each operational day has an associated fixed cost of $\kappa_{0}\geq 0$. Similarly, vehicles and pollsters have fixed hiring costs of $\kappa_{1}\geq 0$ and $\kappa_{2}\geq 0$, respectively, and they work within a daily time frame $[0,B_{\max}]$. Due to labor regulations, pollsters must take a break of $P\geq 0$ minutes starting within the time interval $[T_{0},T_{1}]\subset[0,B_{\max}-P]$.
@@
A vehicular route is a simple path from node 00 to $12n+$ 1 using exclusively arcs in $A_{V}$. A pedestrian route, also called a walking path, is a simple path from $i\in C_{-}$ to $j\in C_{+}$ consisting of an alternating sequence of arcs in the sets $A_{S}$ and $A_{W}$. A feasible service route for a pollster is defined as a simple path from 00 to $12n+$ 1 linking pedestrian routes with segments of vehicular routes. Thus, at every node $i$, a pollster spends $t_{i,i+n}$ units of time to accomplish a task, while respecting the maximum daily duty length. Finally, a feasible daily plan consists of different feasible service routes synced according to the duration of each pedestrian and vehicular route, and any sequence of feasible daily plans forms an aggregated daily plan.
@@
The Integrated Vehicle and Pollster Routing Problem (IVPRP) consists of finding an aggregated daily plan to visit every node $i\in C$ exactly once at minimum cost. This plan must hire at most $|K|$ vehicles with capacity $Q$ and $|E|$ pollsters within no more than $|S|$ days. The IVPRP is $\mathcal{NP}$–hard, as proved in the following theorem.
@@
The $\mathcal{NP}$–hardness is proved by using a polynomial transformation from the Traveling Salesman Problem (TSP). Let $K_{n+1}$ ($n\geq 3$) be a complete graph with a set of nodes $\{0,1,\dots,n\}$ and weights $c:E(K_{n+1})\to\mathbb{R}_{+}$. An instance of the IVPRP is constructed as follows: an arbitrary node in $V(K_{n+1})$ is chosen as the depot. Without loss of generality, let 00 be such a node. The remaining nodes in $V(K_{n+1})$ correspond to every store in our problem, i.e., $C\coloneqq C_{-}\cup C_{+}$, where $C_{-}=V(K_{n+1})\setminus\{0\}$. The multigraph $D$ is constructed as described above. The cost function $c$ in the TSP corresponds to the vehicular travel times $\tau$ for arcs $(i,j)$, $i,j\in C$, $i\neq j$ in $A_{V}$, and 00 otherwise. Pedestrian travel time is set to $t_{a}\gg\sum_{a\in A_{V}}\tau_{a}$, and $t_{a}=0$ for $a\in A_{S}$. Moreover, fix $|E|=|K|=|S|=1$, $Q=1$, $P=0$, $B_{\max}=\infty$, and the break time window is set to $[0,B_{\max})$.
If the Hamiltonian cycle $\{0,i,j,k,\dots,0\}$ is a feasible solution of the TSP with cost $H$, then the IVPRP has a feasible solution with the same cost. This solution is given by the vehicular route $\{0,i,i+n,j,j+n,k,k+n,\dots,2n+1\}$, the pedestrian route $\bigcup_{i\in C_{-}}(i,i+n)$, and both forming a service route, which is also a feasible daily plan.
Conversely, suppose that the IVPRP has a feasible solution with an aggregated daily plan of length less than $\sum_{a}\tau_{a}$. Since the pedestrian travel times are larger than $\sum_{a}\tau_{a}$, the pedestrian arcs in $A_{W}$ are never used in that solution. Thus, the feasible service route reveals a Hamiltonian cycle for the TSP with the same cost. ∎
The $\mathcal{NP}$–hardness condition for the Integrated Vehicle and Pollster Routing Problem makes it unlikely to find an optimal solution in polynomial time. As a consequence, Section 3 is devoted to identifying heuristic methods to obtain good-quality solutions in a reasonable amount of time.